OpenStudy (anonymous):

Solve, using the substitution method. y = 3x + 5 4x – y = 5 (10, 35) (15, 10) There are an infinite number of solutions. There is no solution.

OpenStudy (mikeyymonroe):

just plug in the y = 3x + 5 into the 4x - y = 5 and solve then do the for the other equation.

OpenStudy (anonymous):

i said it d

OpenStudy (mikeyymonroe):

hb you work out the problem before choosing an answer hun k?

OpenStudy (anonymous):

i did do it and theres nothing so there fore picked D okay hun

OpenStudy (mikeyymonroe):

as long as you did your work, :)

OpenStudy (anonymous):

it is D ? or not

OpenStudy (anonymous):

no its C

OpenStudy (mikeyymonroe):

message me how you did your work and ill tell you the correct answer.

OpenStudy (anonymous):

all i did was input them in and it didnt go with it so yaa thats how i did it why cant u just tell me whats the big deal

OpenStudy (mikeyymonroe):

ok the answer is A

OpenStudy (anonymous):

no its not

OpenStudy (mikeyymonroe):

when you subsitute the y= you get 4x - 3x + 5 = 5 nd you simplify and you get x + 5 = 5 so x=10 then you plug that into y = 3x + 5 and you get y = 3(10) + 5 and you simplify and get y = 35. x=10, y=35

OpenStudy (mikeyymonroe):

answer is A

OpenStudy (mikeyymonroe):

do you understand how i got that?

OpenStudy (anonymous):

yes i did

OpenStudy (mikeyymonroe):

ok :)